# SPARSE: Scattering Poles and Amplitudes from Radial Schrödinger Equations

### v1

@@ Abstract
We introduce an algorithm for the solution of a three-dimensional, two-body Schrödinger equation for scattering states. The equation is first reduced to a system of coupled radial Schrödinger equations. The system of differential equations is approximated as an ordinary linear nonhomogeneous system using the finite difference method. Dirichlet boundary conditions are imposed at the origin and at an arbitrary large radius. The physical $K$-matrix for real energies is calculated from the numerical solutions of this system by comparison to the analytical real solutions for large distances. Scattering amplitudes are calculated from the physical $K$-matrix. Scattering poles are obtained by extrapolating the $K$-matrix to complex energies.
PROGRAM SUMMARY
Program Title: SPARSE
CPC Library link to program files: (to be added by Technical Editor)
Developer’s repository link: github.com/Robrusch/SPARSE
Licensing provisions: MIT
Programming language: Python
Nature of problem: Calculation of two-body scattering poles and amplitudes from a radial multichannel Schrödinger equation.
Solution method: We approximate the second derivative using the finite difference method. We impose Dirichelet boundary conditions and reduce the differential equation to a nonhomogeneous linear system. We obtain the $K$-matrix from integrals of the wavefunctions times a sine or cosine function calculated using the trapezoidal rule. We obtain the scattering amplitudes from the $K$-matrix for real energies. We obtain the scattering poles by extrapolating the $K$-matrix to complex energies through a rational approximation.
journal: Computer Physics Communications
@@ 1 Introduction
The Schrödinger equation applies to scattering states just as well as to bound states. In fact, the Schrödinger equation provides the foundation of many approximate analytical methods to calculate scattering resonances using nonrelativistic quantum mechanics. But the advantages of calculating scattering poles and amplitudes directly from the Schrödinger equation are often overlooked.
@@
The most obvious benefit of calculating scattering states directly from the Schrödinger equation is that it does not require any additional assumption, approximation, or parametrization that may undermine the validity of the results. The scattering amplitudes are completely determined by the Schrödinger equation, and thus they are just as good as the underlying potential. That is most advantageous for those cases where resonant line shapes overlap with one another or with the energy thresholds where scattering channels open up. There is no need of a special treatment for esoteric structures like dips and cusps in the scattering amplitudes, since they are automatically included in the Schrödinger equation.
@@
A multichannel scattering Schrd̈inger equation can be used to describe the quantum mechanical scattering of particles in systems where different final states are possible after the collision. One notable example is nucleon-nucleon scattering, where the state of the colliding particles can change at relatively low scattering momentum ($NN\to NN,N\Delta,NN^{\prime},\Delta\Delta,\dots$).
@@
In this paper, we introduce a simple yet effective algorithm for calculating scattering poles and amplitudes directly from a multichannel Schrödinger equation. Our method is to solve the Schrödinger equation using a simple finite-difference method (see Chapter 3 of Reference [1]). Other common methods include propagation algorithms [2, 3] or basis expansions [4, 5] (see Reference [1] for a review).
@@
The rest of the paper is structured as follows. In Section 2, we derive the multichannel radial Schrödinger equation describing an elastic or inelastic 2-body scattering process. In Section 3, we reduce the radial Schrödinger equation down to a nonhomogeneous linear system. In Section 4, we calculate the reactance matrix $K$ for the scattering process by comparing the numerical solutions of the Schrödinger equation to the analytic ones for scattering states with real boundary conditions. In Section 5, we review the calculation of scattering amplitudes from the $K$-matrix and we calculate scattering resonances by extrapolating to complex energies using a rational approximation. In Section 6, we showcase the SPARSE algorithm using a relatively simple example. Finally, we summarize these results in Section 7.
@@ 2 Radial Schrödinger equation
A two-body elastic or inelastic scattering process can be described by a multichannel Schrödinger equation whose most general form is
@@
where $\bm{r}_{1}$, $\bm{r}_{2}$ are the particles’ coordinates , $\bm{p}_{1}=-i\bm{\nabla}_{\bm{r}_{1}}$, $\bm{p}_{2}=-i\bm{\nabla}_{\bm{r}_{2}}$ are their conjugate momenta, $m_{1}$ and $m_{2}$ are diagonal mass matrices, $V(\bm{r}_{1}-\bm{r}_{2})$ is a spin-dependent potential matrix, $E$ is the total energy, and $\psi(\bm{r}_{1},\bm{r}_{2})$ is a multichannel wavefunction with spin. Note that we use natural units, $\hbar=c=1$.
We wish to reduce Equation (1) down to a differential equation in a single vector variable. Let us introduce the relative position
@@
which are the momenta conjugate to $\bm{r}$ and $\bm{R}$, $\bm{p}=-i\bm{\nabla}_{\bm{r}}$ and $\bm{P}=-i\bm{\nabla}_{\bm{R}}$. Expressing $\bm{r}_{1}$, $\bm{r}_{2}$, $\bm{p}_{1}$, and $\bm{p}_{2}$ inside Equation (1) in terms of $\bm{r}$, $\bm{R}$, $\bm{p}$, and $\bm{P}$ using Equations (2)-(5), we obtain
@@
is a diagonal reduced-mass matrix. The Schrödinger equation is separable if the matrix multiplying $\bm{p}\cdot\bm{P}$ is zero, which requires
@@
Equation (10) is a differential equation for a multichannel vector wavefunction $\psi(\bm{r})$ of the relative position vector $\bm{r}$. We wish to further reduce it to an equation for a multichannel radial wavefunction $\psi(r)$ of the distance $r=\lvert\bm{r}\rvert$. Let us label the channels of the Schrödinger equation by $a$. The Schrödinger equation in the channel $a$ reads
@@
where $\psi_{b}(\bm{r})$ is a vector wavefunction component and $V_{ab}(\bm{r})$ are potential matrix elements. Let $\alpha_{a}$ and $\beta_{a}$ be the two particles in the channel $a$ with spins $J_{\alpha_{a}}$ and $J_{\beta_{a}}$, respectively. We can expand the vector wavefunction $\psi_{a}(\bm{r})$ as
@@
where $\psi_{a,l,s}^{\sigma,m}(r)$ are radial wavefunctions, $Y_{l}^{m}(\theta,\phi)$ are spherical harmonics, and $\xi_{s}^{\sigma}$ are irreducible representations from the decomposition of the direct product of the spin states $\chi_{J_{\alpha_{a}}}^{M_{\alpha_{a}}}$ and $\chi_{J_{\beta_{a}}}^{M_{\beta_{a}}}$ of the two particles $\alpha_{a}$ and $\beta_{a}$,
$\xi_{s}^{\sigma}=\sum_{M_{\alpha_{a}},M_{\beta_{a}}}\langle J_{\alpha_{a}}J_{\beta_{a}}M_{\alpha_{a}}M_{\beta_{a}}|J_{\alpha_{a}}J_{\beta_{a}}s\sigma\rangle\chi_{J_{\alpha_{a}}}^{M_{\alpha_{a}}}\chi_{J_{\beta_{a}}}^{M_{\beta_{a}}},$ (14)
where the term with angle brackets is a Clebsch-Gordan coefficient. The sum over the spin quantum number $s$ ranges from $\lvert J_{\alpha_{a}}-J_{\beta_{a}}\rvert$ to $J_{\alpha_{a}}+J_{\beta_{a}}$, and for each $s$ the sum over the third component $\sigma$ ranges from $-s$ to $+s$. The sum over the orbital-angular-momentum quantum number $l$ ranges from 0 to infinity, and for each $l$ the sum over the third component $m$ ranges from $-l$ to $+l$. The radial wavefunctions in Equation (13) are labeled by quantum numbers $\sigma$ and $m$ for the third components of the spin and orbital angular momentum. It is more convenient to rewrite Equation (13) as an expansion in radial wavefunctions labeled by quantum numbers $J$ and $M$ for the total angular momentum,
@@
The sum over $J$ ranges from either 0 or $\frac{1}{2}$ to infinity, and for each $J$ the sum over the third component $M$ ranges from $-J$ to $+J$.
We derive the Schrödinger equation for the radial wavefunctions through the following steps:
@@
    Multiply on the left by $\sum_{\sigma^{\prime},m^{\prime}}\langle s^{\prime}l^{\prime}\sigma^{\prime}m^{\prime}|s^{\prime}l^{\prime}J^{\prime}M^{\prime}\rangle Y_{l^{\prime}}^{m^{\prime}\ast}(\theta,\phi)\xi_{s^{\prime}}^{\sigma^{\prime}{\dagger}}$.
@@
    Integrate over the solid angle $\Omega$.
@@
Once the above procedure has been applied to each channel $a$, the Schrödinger equation for the multichannel radial wavefunction $\psi^{JM}(r)$ with total angular momentum $J$ and third component $M$ can be written concisely as
$\biggl{[}\frac{1}{2\mu}\biggl{(}-\frac{\mathrm{d}^{2}\hphantom{r}}{\mathrm{d}r^{2}}-\frac{2}{r}\frac{\mathrm{d}\hphantom{r}}{\mathrm{d}r}+\frac{L(L+1)}{r^{2}}\biggr{)}+V^{J}(r)-E\biggr{]}\psi^{JM}(r)=0,$ (16)
@@
where $L$ is a diagonal orbital-angular-momentum matrix, $V^{J}(r)$ is a radial potential matrix, and each channel $a$ of the three-dimensional Schrödinger equation has been expanded into spin-orbital channels $(a,l,s)$. Note that $s$ and $l$ are constrained by the triangle relations $\lvert J_{\alpha_{a}}-J_{\beta_{a}}\rvert\leq s\leq J_{\alpha_{a}}+J_{\beta_{a}}$ and $\lvert J-s\rvert\leq l\leq J+s$, so that each channel $a$ expands into a finite number of spin-orbital channels $(a,l,s)$ for each $J$. To simplify the notation, we relabel the spin-orbital channels of the radial wavefunction by a single integer $i$. Then the radial potential matrix element for the spin-orbital channels $i$ and $j$ is given by
@@
$V_{ij}^{J}(r)=\sum_{\sigma_{i},m_{i}}\sum_{\sigma_{j},m_{j}}\langle s_{i}l_{i}\sigma_{i}m_{i}|s_{i}l_{i}JM\rangle\langle s_{j}l_{j}\sigma_{j}m_{j}|s_{j}l_{j}JM\rangle\\ \times\int\mathrm{d}\Omega\,Y_{l_{i}}^{m_{i}\ast}(\theta,\phi)Y_{l_{j}}^{m_{j}}(\theta,\phi)\xi_{s_{i}}^{\sigma_{i}{\dagger}}V_{ab}(\bm{r})\xi_{s_{j}}^{\sigma_{j}},$ (17)
@@
where $(a,s_{i},l_{i})\mapsto i$, $(b,s_{j},l_{j})\mapsto j$, and the spin vectors $\xi_{s_{i}}^{\sigma_{i}{\dagger}}$ and $\xi_{s_{j}}^{\sigma_{j}}$ can be expanded in terms of the spin vectors of the individual particles $\chi_{J_{\alpha_{i}}}^{M_{\alpha_{i}}},\chi_{J_{\beta_{i}}}^{M_{\beta_{i}}}$ and $\chi_{J_{\alpha_{j}}}^{M_{\alpha_{j}}},\chi_{J_{\beta_{j}}}^{M_{\beta_{j}}}$ using Equation (14). Note that the right side of Equation (17) does not actually depend on $M$ by virtue of the Wigner-Eckart theorem.
@@
There is a somewhat common misunderstanding that the reduction of a three-dimensional Schrödinger equation to a radial form can be performed only for a spherically-symmetric potential that does not depend on the angles $\theta$ and $\phi$. However, in our derivation of Equations (16) and (17) we have made no assumption on the potential $V(\bm{r})$ other than it conserves angular momentum. In the trivial case where the potential is independent of spin, then conservation of angular momentum does indeed require that it is also spherically symmetric. If the potential does depend on spin, then it can also depend on the angles $\theta$ and $\phi$ but the reduction to Equation (16) would still apply. A brute-force method to determine if any given potential $V(\bm{r})$ conserves angular momentum is to evaluate the right side of Equation (17) with $J$ and $M$ in the first Clebsch-Gordan coefficient substituted by $J^{\prime}$ and $M^{\prime}$. The potential $V(\bm{r})$ conserves angular momentum if and only if the result is zero unless $J^{\prime}=J$ and $M^{\prime}=M$ and it is independent of $M$.
@@
Equation (20) is the form of the Schrödinger equation solved by the SPARSE algorithm. It has $N$ spin-orbital channels labeled by an integer $i=1,2,\dots,N$. The channel $i$ has spin $s_{i}$ and orbital angular momentum $l_{i}$. The reduced-mass matrix elements are $\mu_{ij}=\delta_{ij}\mu_{i}$. The orbital-angular-momentum matrix elements are $L_{ij}=\delta_{ij}l_{i}$. The radial potential matrix elements $V_{ij}(r)$ can be determined by evaluating Equation (17) for a given potential $V(\bm{r})$.

### v2

@@ Abstract
We introduce an algorithm for the solution of a large system of radial Schrödinger equations for scattering states. The system of differential equations is approximated as an ordinary linear nonhomogeneous system using the finite difference method. Dirichlet boundary conditions are imposed at the origin and at an arbitrary large radius. The $K$-matrix for physical energies is calculated from the numerical solutions of the system by comparison to the analytical real solutions at large distances. Scattering poles and amplitudes are calculated from the physical $K$-matrix.
GitHub repository link: github.com/Robrusch/SPARSE
@@ 1 Introduction
The Schrödinger equation applies to scattering states just as well as to bound states. In fact, the Schrödinger equation provides the foundation of many approximate analytical methods to calculate scattering poles and amplitudes using nonrelativistic quantum mechanics. But the advantages of calculating scattering poles and amplitudes directly from the Schrödinger equation are often overlooked.
@@
The most obvious benefit of calculating scattering states directly from the Schrödinger equation is that it does not require any secondary approximation that may deteriorate the accuracy of the results. The scattering wavefunctions are completely determined by the Schrödinger equation, which in turn is completely determined by the potential and the masses of the degrees of freedom. Therefore, the scattering poles and amplitudes obtained in this way are just as good as the model or framework in which the Schrödinger equation is defined. That is most advantageous for those cases where resonant line shapes overlap with one another or with the energy thresholds where scattering channels open up. There is no need for a special treatment of esoteric structures like dips and cusps in the scattering amplitudes, since they are automatically included in the Schrödinger equation.
@@
A system of Schrödinger equations can be used to describe the quantum mechanical scattering of particles in systems where different final states are possible after a collision at nonrelativistic center-of-mass energies. One notable example is nucleon-nucleon scattering, where inelastic effects start to play a role at relatively low scattering momenta ($NN\to NN,N\Delta,NN^{\prime},\Delta\Delta,\dots$).
@@
In this paper, we introduce a simple yet effective algorithm for calculating scattering poles and amplitudes directly from a system of radial Schrödinger equations. Our method is to solve the system using a simple finite-difference method (see Chapter 3 of Reference [1]). Other common methods include propagation algorithms [2, 3] or basis expansions [4, 5] (see Reference [1] for a review). The crudeness of the finite-difference method is offset by its efficiency, which makes it the gold standard for numerically solving a differential equation with minimal effort.
In phenomenological applications, there are many situations in which one has to solve a system of Schrödinger equations several times to obtain a definite prediction. For instance, one may need to fit the parameters of some potential model, or one may want to estimate the error on the calculated spectrum from the uncertainty of the potential parameters using a bootstrap method. In those cases, it is desirable to have an algorithm that solves the system of Scrhödinger equations as efficiently as possible. The extreme efficiency of the finite-difference method allows to solve systems with tens of coupled Schrödinger equations [^2] Such as the equations governing certain strongly interacting systems in the Born-Oppenheimer approximation for quantum chromodynamics; see, for instance, Ref. [6]. using only modest computing power and time. That is the main purpose and novelty of the SPARSE algorithm.
@@
The rest of the paper is structured as follows. In Section 2, we derive the system of radial Schrödinger equations describing an inelastic two-body scattering process in a partial wave with definite total-angular-momentum quantum numbers. In Section 3, we reduce the system of radial Schrödinger equations down to a nonhomogeneous linear system. In Section 4, we calculate the reactance matrix $K$ for the scattering process by comparing the numerical wavefunctions to the analytic ones for scattering states with real boundary conditions. In Section 5, we review the calculation of scattering poles and amplitudes from the $K$-matrix. In Section 6, we showcase the SPARSE algorithm using a simple example that serves also as a tutorial. Finally, we summarize these results in Section 7.
@@ 2 Radial Schrödinger equation
A two-body elastic or inelastic scattering process at nonrelativistic energies can be described by a system of coupled Schrödinger equations. Using matrix notation, the system can be written concisely as
@@
where $\bm{r}_{1}$, $\bm{r}_{2}$ are the particles’ coordinates , $\bm{p}_{1}=-i\bm{\nabla}_{\bm{r}_{1}}$, $\bm{p}_{2}=-i\bm{\nabla}_{\bm{r}_{2}}$ are their conjugate momenta, $m_{1}$ and $m_{2}$ are diagonal mass matrices, $V(\bm{r}_{1}-\bm{r}_{2})$ is a spin-dependent potential matrix, $E$ is the total energy, and $\psi(\bm{r}_{1},\bm{r}_{2})$ is a vector wavefunction. Note that we use natural units, $\hbar=c=1$. Each Schrödinger equation in the system governs an elastic scattering process in a given channel. Each diagonal element of the mass matrices $m_{1}$ and $m_{2}$ contains the mass of the particles 1 and 2 in the corresponding channel. The couplings between the different Schrödinger equations introduce inelastic effects. In general, Equation (1) describes interacting particles with spin. Each component of the vector wavefunction $\psi(\bm{r}_{1},\bm{r}_{2})$ contains the spin states of the scattering particles in the corresponding channel. Each element of the potential matrix $V(\bm{r}_{1}-\bm{r}_{2})$ is an operator acting on the spin states.
We wish to reduce Equation (1) down to a system of differential equations in a single vector variable. Let us introduce the relative position
@@
which are the momenta conjugate to $\bm{r}$ and $\bm{R}$, that is, $\bm{p}=-i\bm{\nabla}_{\bm{r}}$ and $\bm{P}=-i\bm{\nabla}_{\bm{R}}$. Expressing $\bm{r}_{1}$, $\bm{r}_{2}$, $\bm{p}_{1}$, and $\bm{p}_{2}$ inside Equation (1) in terms of $\bm{r}$, $\bm{R}$, $\bm{p}$, and $\bm{P}$ using Equations (2)-(5), we obtain
@@
is a diagonal reduced-mass matrix. The system of Schrödinger equations is separable if the matrix multiplying $\bm{p}\cdot\bm{P}$ is zero, which requires
@@
We now wish to further reduce Equation (10) to a system of differential equations in the distance $r=\lvert\bm{r}\rvert$. Let us label the different Schrödinger equations in the system by an integer $a$. The Schrödinger equation in channel $a$ reads
@@
where $\psi_{a}(\bm{r})$ is the wavefunction in channel $a$ and $V_{ab}(\bm{r})$ are potential matrix elements. Recall that $\psi_{a}(\bm{r})$ contains the spin states of the scattering particles and $V_{ab}(\bm{r})$ is an operator acting on the spin states. Let $J_{1,a}$ and $J_{2,a}$ be the spins of the particles 1 and 2 in the channel $a$. We can expand the wavefunction $\psi_{a}(\bm{r})$ as
@@
where $\psi_{a,s,l}^{\sigma,m}(r)$ are radial wavefunctions, $Y_{l}^{m}(\theta,\phi)$ are spherical harmonics, and $\xi_{s}^{\sigma}$ are irreducible representations from the decomposition of the direct product of the spin states $\chi_{J_{1,a}}^{M_{1,a}}$ and $\chi_{J_{2,a}}^{M_{2,a}}$ of the particles 1 and 2,
$\xi_{s}^{\sigma}=\sum_{M_{1,a},M_{2,a}}\langle J_{1,a}J_{2,a}M_{1,a}M_{2,a}|J_{1,a}J_{2,a}s\,\sigma\rangle\chi_{J_{1,a}}^{M_{1,a}}\chi_{J_{2,a}}^{M_{2,a}},$ (14)
where the term with angle brackets is a Clebsch-Gordan coefficient. The sum over the total spin quantum number $s$ ranges from $\lvert J_{1,a}-J_{2,a}\rvert$ to $J_{1,a}+J_{2,a}$, and for each $s$ the sum over the third component $\sigma$ ranges from s-s to +s+s. The sum over the orbital-angular-momentum quantum number $l$ ranges from 0 to infinity, and for each $l$ the sum over the third component $m$ ranges from l-l to +l+l. The radial wavefunctions in Equation (13) are labeled by quantum numbers $\sigma$ and $m$ for the third components of the spin and orbital angular momentum, respectively. It is more convenient to rewrite Equation (13) as an expansion in radial wavefunctions labeled by quantum numbers $J$ and $M$ for the total angular momentum,
@@
The sum over $J$ ranges from either 0 or $\frac{1}{2}$ to infinity, and for each $J$ the sum over the third component $M$ ranges from J-J to +J+J.
We derive the system of radial Schrödinger equations through the following steps:
@@
    Multiply on the left by $\sum_{\sigma^{\prime},m^{\prime}}\langle s^{\prime}l^{\prime}\sigma^{\prime}m^{\prime}|s^{\prime}l^{\prime}J^{\prime}M^{\prime}\rangle Y_{l^{\prime}}^{m^{\prime}}(\theta,\phi)^{\ast}\xi_{s^{\prime}}^{\sigma^{\prime}{\dagger}}$.
@@
    Integrate over the angles $\theta$ and $\phi$.
@@
Once the above procedure has been applied to each channel $a$, one obtains infinitely many systems of radial Schrödinger equations labeled by conserved quantum numbers $J$ and $M$ for the total angular momentum and its third component. Using matrix notation, the system for given quantum numbers $J$ and $M$ can be written concisely as
$\biggl[\frac{1}{2\mu}\biggl(-\frac{\mathrm{d}^{2}\hphantom{r}}{\mathrm{d}r^{2}}-\frac{2}{r}\frac{\mathrm{d}\hphantom{r}}{\mathrm{d}r}+\frac{L(L+1)}{r^{2}}\biggr)+V^{J}(r)-E^{J}\biggr]\psi^{JM}(r)=0,$ (16)
@@
where $L$ is a diagonal orbital-angular-momentum matrix, $V^{J}(r)$ is a radial potential matrix, and $\psi^{JM}(r)$ is a vector wavefunction. Note that the wavefunctions depend on both $J$ and $M$, but the potential and the energy depend only on $J$. The radial Schrödinger equations in this system are labeled by triplets $(a,l,s)$ that identify spin-orbital channels of the scattering process. Since $s$ and $l$ are constrained by the triangle relations $\lvert J_{1,a}-J_{2,a}\rvert\leq s\leq J_{1,a}+J_{2,a}$ and $\lvert J-s\rvert\leq l\leq J+s$, the system in Equation (16) contains only a finite number of radial Schrödinger equations for any value of $J$. To simplify the notation, we replace the labels $(a,l,s)$ for the spin-orbital channels by a single integer $i$. Then the radial potential matrix elements can be expressed as
@@
$V_{ij}^{J}(r)=\sum_{\sigma_{i},m_{i}}\sum_{\sigma_{j},m_{j}}\langle s_{i}l_{i}\sigma_{i}m_{i}|s_{i}l_{i}JM\rangle\langle s_{j}l_{j}\sigma_{j}m_{j}|s_{j}l_{j}JM\rangle\\ \times\int\mathrm{d}\Omega\,Y_{l_{i}}^{m_{i}}(\theta,\phi)^{\ast}Y_{l_{j}}^{m_{j}}(\theta,\phi)\xi_{s_{i}}^{\sigma_{i}{\dagger}}V_{ab}(\bm{r})\xi_{s_{j}}^{\sigma_{j}},$ (17)
@@
where $(a,s_{i},l_{i})\to i$, $(b,s_{j},l_{j})\to j$, and the total spin states $\xi_{s_{i}}^{\sigma_{i}}$ and $\xi_{s_{j}}^{\sigma_{j}}$ can be expanded in terms of the spin states $\chi_{J_{1,i}}^{M_{1,i}},\chi_{J_{2,i}}^{M_{2,i}}$ and $\chi_{J_{1,j}}^{M_{1,j}},\chi_{J_{2,j}}^{M_{2,j}}$ of particles 1, 2 in the channel $i$ and $j$ using Equation (14). Note that the right side of Equation (17) does not depend on $M$ by virtue of the Wigner-Eckart theorem.
@@
There is a somewhat common misunderstanding that the reduction of a three-dimensional Schrödinger equation to a radial form can be performed only for a spherically-symmetric potential that does not depend on the angles $\theta$ and $\phi$. However, in our derivation of Equations (16) and (17) we have made no assumption on the potential $V(\bm{r})$ other than it conserves angular momentum. In the trivial case where the potential is independent of spin, then conservation of angular momentum does indeed require that it is also spherically symmetric. If the potential does depend on spin, then it can also depend on the angles $\theta$ and $\phi$ but Equations (16) and (17) would still be valid. A brute-force method to determine if any given potential $V(\bm{r})$ conserves angular momentum is to evaluate the right side of Equation (17) with $J$ and $M$ in one of the two Clebsch-Gordan coefficients replaced by $J^{\prime}$ and $M^{\prime}$. The potential $V(\bm{r})$ conserves angular momentum if and only if the result is zero for $J^{\prime}\neq J$ or $M^{\prime}\neq M$ and it is independent of $M$ when $J^{\prime}=J$ and $M^{\prime}=M$.
@@
Equation (20) is the system of radial Schrödinger equations solved by the SPARSE algorithm. It has $N$ spin-orbital channels labeled by an integer $i=1,2,\dots,N$. The channel $i$ has spin $s_{i}$ and orbital angular momentum $l_{i}$. The reduced-mass matrix elements are $\mu_{ij}=\delta_{ij}\mu_{i}$. The orbital-angular-momentum matrix elements are $L_{ij}=\delta_{ij}l_{i}$. The radial potential matrix elements $V_{ij}(r)$ are given in Equation (17).